# RTLCoder: Fully Open-Source and Efficient LLM-Assisted RTL Code Generation Technique

### v4

@@ III New Training Scheme Incorporating Code Quality Feedback
Besides the new training dataset, we propose a new LLM training scheme that incorporates code quality scoring. It significantly improves the RTLCoder’s performance on the RTL generation task. Also, we revised the training process from the algorithm perspective to reduce the GPU memory consumption of this new training method, allowing implementation with limited hardware resources.
III-A Existing Supervised Training on LLMs
Figure 13: Comparison between (a) existing MLE-based LLM training flow and (b) our proposed LLM training flow.
This part will first introduce the existing supervised training method for LLMs. Then we will further discuss its limitations in RTL generation tasks. Suppose we have a training data dateset $\left\{x_{i},y_{i}\right\}$ for $i=1,...,N$, where $x_{i}$ represents an design instruction, $y_{i}$ represents the corresponding correct reference code. Each sample of data will be split into a sequence of tokens by certain rules during the preprocessing process. In this paper, we use $x_{i}=\left\{x_{i}^{t}\right\}$ and $y_{i}=\left\{y_{i}^{t}\right\}$ for $t=1,2,...,T$ to represent the tokenized sequence.
LLMs generate a sequence by continuously predicting the next token based on the already generated previous ones. For a decoder-only language model, which is the mainstream LLM architecture, the probability of producing the next token depends only on the previous output tokens and the input instruction. We denote the probability of generating the $t$-th token $r_{t}$ ($r_{t}$ can be any single token in the vocabulary) as $P_{\pi}\left(r_{t}\mid x_{i},y_{i}^{<t}\right)$ where $\pi$ represents the model parameters and $y_{i}^{<t}$ denotes the already generated previous tokens $\left\{y_{i}^{1},..,y_{i}^{t-1}\right\}$. Then the log probability of generating the whole sequence can be written as: $\sum_{t=1}^{T}\log P_{\pi}\left(y_{i}^{t}\mid x_{i,},y_{i}^{<t}\right)$.
In the existing training method, Maximum Likelihood Estimation (MLE) is commonly used to find the best parameters $\pi$ that maximize the log probability. The training flow is shown in Figure 13(a). The loss is usually defined as below:
$loss_{mle}=-\sum_{t=1}^{T}\log P_{\pi}\left(y_{i}^{t}\mid x_{i,},y_{i}^{<t}\right)\vspace{-.03in}$
@@ III-A Existing Supervised Training on LLMs
However, there exists a phenomenon named exposure bias [26, 27]. Since the above sequence generation is autoregressive, which means the model always predicts the next token based on its own generated previous ones $r_{i}^{<t}$ rather than the reference tokens $y_{i}^{<t}$. Therefore, even though the probability of producing $y_{i}^{t}$ is high when given $y_{i}^{<t}$ in the training, it can still result in a huge deviation from the reference code in the generation process.
We have also observed this phenomenon in our experiments. After the supervised training, the qualities of multiple generated code candidates for the same instruction may diverse greatly in the performance aspect. They can include correct code while at the same time including many low-quality answers. Some candidates exhibit serious nonsense duplication [^8] We notice that this duplication couldn’t be simply dealt with by adding repetition penalty to the decoding process like other works in natural text generation. Because some correct RTL design code also contain similarly repetitive expressions..
To alleviate the exposure bias phenomenon, we suggest that in addition to the reference code $y_{i}$, the model’s generation should also be considered in the training process. Since the generation may be different from the reference code, it is necessary to introduce a scoring mechanism to judge the quality of generated candidates. We will give our detailed solution in Section III-B.
III-B Our Proposed Training Method
@@
Our proposed training scheme is illustrated in Figure 13(b). For each instruction, we will now collect multiple code candidates generated by the initial pre-trained model. Then, we pack these candidates and the original reference code $y_{i}$ together as $\textbf{y}_{i}=\left\{y_{i,k}\right\}$, $k=1,2,..,K$, where $K$ represents the number of generated code for one instruction. Next, all these candidates will be scored by the scoring mechanism $R(x_{i},y_{i,k})$ which could be a syntax checker or unit test for functionality check. We will then obtain a set of score $\textbf{z}_{i}=\left\{z_{i,k}\right\}$, $k=1,2,..,K$, denoting the quality for the code sample $\{y_{i,k}\}$. In the training process, we aim to make the model learn to assign relatively higher generation probabilities to answers with higher scores. In this way, the model not only learns from the reference code, but also from the new information introduced by the quality score feedback.
Algorithm 1 Training scheme using gradients splitting
Input: The single data sample $\left\{x_{i},\textbf{y}_{i},\textbf{z}_{i}\right\}$. Model forward function $s_{i,k}=f_{\pi}(x_{i},y_{i,k},z_{i,k})$. Loss calculation function $L_{\pi}(\textbf{s}_{i},\textbf{z}_{i})$. GPU affordable batch size $J$. Model parameters $w$.
Output: The derivative of the loss with respect to model parameters: $g_{i}$.
1:Group the sample $\left\{x_{i},y_{i,k}\right\}$ for $k=1,2,...,K$ into $Q$ parts based on batch size $J$.
2:initialize empty vector list $temp$. Initialize the gradients $g_{i}=0$.
3:for $q\in Q$ do
4:    Calculate $s_{i,k}=f_{\pi}(x_{i},y_{i,k},z_{i,k})$, for $k\in q$.
5:    Empty the computation graph
6:Calculate $loss=L_{\pi}(\textbf{s}_{i},\textbf{z}_{i})$ //$\textbf{s}_{i}=\left\{s_{i,k}\right\}$ for $k=1,..,K$
7:Backward process: $temp_{k}={\partial\operatorname{loss}}/{\partial s_{i,k}}$, for $k=1,...,K$
8:for $q\in Q$ do
9:    Calculate $s_{i,k}=f_{\pi}(x_{i},y_{i,k},z_{i,k})$, for $k\in q$
10:    Backward process: $g_{i}=g_{i}+\sum_{k\in q}temp_{k}{\partial s_{i,k}}/{\partial w}$
11:    Empty the computation graph
12:Return $g_{i}$
The conditional log probability (length-normalized) of generating the entire code $y_{i,k}$ is commonly written as:
$p_{i,k}=\frac{\sum_{t}\log P_{\pi}\left(y_{i,k}^{t}\mid x_{i},y_{i,k}^{<t}\right)}{\left\|y_{i,k}\right\|}$
We calculate $p_{i,k}$ for all code candidates $\textbf{y}_{i}=\left\{y_{i,k}\right\}$, $k=1,2,..,K$, then we normalize these $p_{i,k}$ values using a $softmax$ function, defining the probability of each code being selected as:
$s_{i,k}=\frac{e^{p_{i,k}}}{\sum_{\tau=1}^{K}e^{p_{i,\tau}}}\vspace{-.03in}$
This $s_{i,k}$ reflects the model’s tendency to output the $k^{\text{th}}$ code candidate, with higher probabilities indicating a greater likelihood that the model will generate it.
To encourage the model to assign higher probability scores to high-quality code, we can define a new loss function term:
$loss_{compare}=\sum_{z_{i,k}<z_{i,\tau}}max\left(s_{i,k}-s_{i,\tau}+\lambda,0\right)\vspace{-.03in}$
where $\lambda$ is a threshold value.
To provide an intuitive explanation of this loss function term, we provide a simple example. Suppose we have the $i^{\text{th}}$ instruction and only two code candidates with initial selection probability $s_{i,1}$ and $s_{i,2}$ with $s_{i,1}+s_{i,2}=1$ and $s_{i,1}>s_{i,2}$. But the first candidate has a lower quality score, i.e., $z_{i,1}<z_{i,2}$. Then the positive loss would drive model parameters to update until the model assigns a new set of $s_{i,1}^{*}$ and $s_{i,2}^{*}$ so that $s_{i,2}^{*}-s_{i,1}^{*}\geq\lambda$ is satisfied.
It is worth noting that this loss only depends on the relative scores among multiple code candidates, so it can still be used when answer quality cannot be precisely quantified. Finally, We define the total loss as:
$loss=loss_{compare}+loss_{mle}$
III-C Reduced Memory by Splitting Gradients
Directly calculating our new $loss$ function even with 1 batch size would still require forwarding all code candidates in a sample at once to maintain all the activation values. This will lead to the $O(K)$ space complexity and make the GPU memory consumption prohibitively high in many large language model training scenarios.
We propose a gradient-splitting approach for model training based on quality score from an algorithm perspective. It can achieve a $O(1)$ space complexity as illustrated in Algorithm 1. The gradients of $loss$ with respect to $w$ can be computed as below:
$\frac{\partial\operatorname{loss}}{\partial w}=\sum_{k}\frac{\partial\operatorname{loss}}{\partial s_{i,k}}\frac{\partial s_{i,k}}{\partial w}\vspace{-.03in}$
The property of the chain rule indicates that we can decompose the gradient updates into several parts. Assume $J$ is the maximum allowable batch size for GPU consumption. We divide the $K$ candidates into $Q$ groups based on the batch size $J$. Firstly, we pass these groups through the forward function separately and collect the obtained $\textbf{s}_{i}$ values as lines 1-5 illustrate. In the second step, we calculate the loss function and compute the derivative of the loss with respect to $\textbf{s}_{i}$ in lines 6-7, storing the temporary results in vector $temp$. In the third step, we perform the forward operation on the original $Q$ groups again and for each forward operation, the obtained $s_{i,k}$ is multiplied by $temp_{k}$ in a dot product, followed by a backward pass to accumulate the gradient in lines 9-12.
@@ IV-A Evaluation Benchmark and Metric
To evaluate the performance of Verilog code generation, there are two representative benchmarks VerilogEval [12] and RTLLM [11].
The VerilogEval [12] benchmark consists of two parts, EvalMachine and EvalHuman, each including more than 100 RTL design tasks. We follow the original paper [12] and use the widely-adopted $pass@k$ metric in code generation tasks:
$\operatorname{pass@k}=E_{i}\left(1-\frac{C_{n-c_{i}}^{k}}{C_{n}^{k}}\right)$
where $n$ is the total number of trials for each instruction and $c_{i}$ is the number of correct code generations for task $i$. We set $n=20$ in this experiment. If any code in the $k$ trials could pass the test, then this task is considered to be addressed and the $pass@k$ metric reflects the estimated proportion of design tasks that could be solved.
The RTLLM V1.1 [11] benchmark contains 29 RTL design tasks at a larger design scale. We mostly follow the testing method in the original paper [11], but further proposes two slightly different metrics for evaluating syntax correctness, using either Synopsys VCS [28] or Design Compiler [29]. They are denoted as Syn-VCS and Syn-DC, respectively. 1) For the Syn-VCS metric, VCS not only requires the design to comply with the Verilog syntax rules, but also requires that the interface of the design correspond to the testbench, so that the circuit can be simulated. 2) For the Syn-DC metric, DC requires the design to be physically synthesizable. The functionality result is obtained by VCS simulation. We calculate the scores of the design syntax part and design functionality part separately. In both parts, following the original benchmark [11], each task is counted as success as long as any of 5 trials passes the test. This can be interpreted as pass@5 metric.
In the generation process, we set ${top}_{p}=0.95$ and $temperature=\left\{0.2,0.5,0.8\right\}$. For all tested models (i.e., baselines, RTLCoder, and ablation studies), we evaluate all 3 $temperature$ conditions and report the best performance for each model.
IV-B Examine Training Set for Fair Evaluation
To ensure a fair evaluation of our proposed RTLCoder, before training, we explicitly examined the similarity between samples in our proposed training dataset and those test cases in benchmarks [12, 11], then we get rid of our training samples that are similar to test cases during the training process.
To measure the similarity between two text sequences, we employed the Rouge-L metric, which is a widely-used similarity calculation scheme in the LLM domain such as by OpenAI [1]. The Rouge-L score $\in[0,1]$, with values closer to 1 indicating higher similarity between the two sequences. For each instruction-code concatenated sample in the training dataset, we computed its Rouge-L value with all test cases in the benchmarks. In addition, we also separately analyzed the distribution of token counts for instructions and code in the dataset. The resulting statistic is in Figure  14.
From Figure  14 (a), we can see that a sample that consists of one instruction and one code candidate is generally within 2048 token length. So we can set 2048 as the max length in our finetuning. In Figure  14 (b), we observed that the majority of training samples in the dataset have a low overlap compared with the benchmark, with Rouge-L scores $<0.3$. However, there are still a small number of samples with higher similarity. To ensure fair evaluation of the RTLCoder, we get rid of all training samples with Rouge-L values $>0.5$ which counts about 100 samples.
Figure 14: Training dataset analysis. (a) Tokens number distribution of instruction and code part. (b) Similarity measurement between training dataset and two benchmarks based on Rouge-L metric.
@@ IV-B Examine Training Set for Fair Evaluation
RTLCoder RTLCoder-Mistral-4bit 7B * 4bit 59.5 72.2 76.9 33.8 42.3 47.1 86.2 41.4
RTLCoder-DeepSeek-4bit 6.7B * 4bit 56.5 73.2 78.4 37.5 50.5 55.5 93.1 37.9
RTLCoder-Mistral 7B 62.5 72.2 76.6 36.7 45.5 49.2 96.6 48.3
@@
IV-C Model Training
@@
Based on our generated dataset with 27K instruction-code pairs, we choose the latest Mistral-7B-v0.1 [23] and DeepSeek-Coder-6.7b [30] as the basic pre-trained model for finetuning. In all experiments, we opted for the Adam optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, and learning rate $\gamma$ = 1e-5, while abstaining from the use of weight decay. Concurrently, we established a context length of 2048 and a global batch size of 256. We trained the model on only 4 consumer-level RTX 4090 GPUs (24GB each), each of which could only afford $2\times 2048$ context length using DeepSpeed stage-2 [31]. Under the hardware constraint, the training is impossible without the proposed gradient-splitting method.
@@
To implement our proposed training scheme, we first generated 3 code candidates for each instruction using the pre-trained model with the Beam search method. Then we use Pyverilog [32] as the syntax checker to score the code candidates. Specifically, we assigned a full score (i.e., 1) for the reference code from the dataset and those candidates who can pass the syntax check. For those who failed syntax checks, we used the Rouge-L metric to assign the code similarity between the candidate and reference code as its score.
In addition, considering GPU memory consumption is a crucial factor that limits the applicability of LLMs, based on quantization methodologies [33], we further quantize the parameters of the obtained RTLCoder into 4 bits, generating RTLCoder-DeepSeek-4bit and RTLCoder-Mistral-4bit, consuming only 4GB memory.
Figure 15: Visualization of key accuracy comparisons from Table III, selecting pass@1 metric on EvalMachine and EvalHuman of  [12]. RTLCoder includes both RTLCoder-Mistral and RTLCoder-DeepSeek. The baseline models include Thakur et al. [13], ChipNeMo [3] and VerilogEval [12].
(a) EvalMachine syntax
(b) EvalMachine functionality
(c) EvalHuman syntax
(d) EvalHuman functionality
Figure 16: Detailed syntax and functionality results of RTLCoder-Mistral on VerilogEval Benchmark [12], reporting EvalMachine and EvalHuman separately. Each sub-figure has 8 columns, and thus cell at $(i,j)$ represents the $((j-1)\times 8+i)^{\text{th}}$ task. The color of each cell indicates the count of correct cases among 20 trials. EvalMachine contains 143 tasks, so the last 1 cell is empty. EvalHuman contains 156 tasks, so the last 4 cells are empty.
Table IV: Detailed Syntax and Functionality Evaluation Results using sampling generation method in RTLLM V1.1 [11]
Design GPT-3.5 GPT-4 Thakur et al. [13] StarCoder[22] RTLCoder-Mistral-4bit RTLCoder-Mistral
Syn-VCS Syn-DC Func Syn-VCS Syn-DC Func Syn-VCS Syn-DC Func Syn-VCS Syn-DC Func Syn-VCS Syn-DC Func Syn-VCS Syn-DC Func
accu 2 2 5 5 4 4 3 4 5 5 4 4
adder_8bit 3 3 4 4 3 3 2 4 5 5 5 5
adder_16bit 1 0 3 3 3 4 2 3 0 0 - 3 3
adder_32b 0 0 - 2 2 1 0 1 3 1 0 1 0
adder_pipe_64b 5 5 5 5 0 0 - 0 0 - 1 1 3 2
multi_booth_8b 5 2 5 5 3 3 4 3 5 5 5 5
multi_16b 5 0 5 5 3 3 3 4 4 2 5 5
multi_pipe_4b 0 0 - 2 2 1 0 3 1 4 1 2 0
multi_pipe_8b 2 0 5 5 3 1 2 3 0 0 - 2 0
div_8bit 3 1 5 1 0 1 - 3 0 3 1 4 1
div_16bit 4 0 5 4 1 2 1 1 0 0 - 0 0 -
JC_counter 5 5 5 5 3 3 4 5 5 5 5 4
right_shifter 4 4 5 5 0 2 - 3 3 5 5 5 5
synchronizer 5 5 4 4 4 4 5 5 4 4 5 5
counter_12 5 5 5 5 2 4 2 4 5 5 5 5
freq_div 5 5 5 5 4 4 4 4 5 5 5 3
signal_gen 5 5 5 5 4 5 4 4 5 5 5 5
serial2parallel 4 4 5 5 4 4 4 4 5 3 5 3
parallel2serial 2 2 5 5 1 2 3 4 3 3 3 2
pulse_detect 4 4 5 3 4 3 3 3 5 5 2 2
edge_detect 5 5 5 5 4 5 3 4 4 2 5 4
FSM 5 4 5 2 4 4 5 5 4 4 5 5
width_8to16 4 3 5 5 4 1 3 4 5 5 5 4
traffic_light 4 0 4 3 5 2 5 3 4 0 4 3
calendar 5 5 5 5 2 1 5 4 1 0 5 5
RAM 4 0 5 2 5 5 2 0 3 0 3 0
asyn_fifo 0 0 - 3 2 0 0 - 0 0 - 0 2 - 1 3
ALU 2 0 - 5 4 - 2 2 1 0 2 1 1 0
PE 5 5 5 5 3 3 3 5 1 1 5 5
Success rate 89.7% 65.5% 11/29 100% 100% 19/29 86.2% 86.2% 7/29 93.1% 82.8% 8/29 86.2% 75.9% 12/29 96.6% 79.3% 14/29
IV-D Experiment Results Overview
@@
Table III summarizes the comparison of all relevant RTL generation solutions, including commercial models GPT3.5/GPT4, models customized for Verilog generation [12, 13, 19], software code generators [21, 22, 23, 30], our proposed RTLCoder and quantized version RTLCoder-4bit, and ablation studies of RTLCoder. In addition, we further visualize key results on VerilogEval benchmark in Figure 15.
@@
In the VerilogEval benchmark [12], for both EvalHuman and EvalMachine categories, RTLCoder-DeepSeek scores 61.2 and 41.6 respectively. It clearly outperforms GPT-3.5 and is only inferior to GPT-4 among all the models in EvalHuman. Specifically, in the EvalMachine part, RTLCoder-DeepSeek and RTLCoder-Mistral even outperforms GPT4 by an absolute value of 1.2% and 2.5%. A similar trend can be observed in the RTLLM benchmark V1.1 [11]. RTLCoder is also second only to GPT-4. In summary, RTLCoder outperforms GPT-3.5 and all non-commercial baseline models in all metrics on both benchmarks. It is surprising that the lightweight RTLCoder with only 7 billion parameters could achieve such impressive accuracy despite its smaller size.
@@
Furthermore, we validate the effectiveness of our proposed dataset and algorithm through an ablation study. The RTLCoder-Mistral-Direct and RTLCoder-DeepSeek-Direct are directly trained with the existing method mentioned in Figure 13(a). Using our training dataset, they can already significantly outperform the base model and even GPT-3.5 on part of these indexes. Then the RTLCoders trained with our proposed training scheme further outperform those using Direct training method on all benchmarks, indicating that our training method greatly further improves the model performance.
In addition, although the quantized model RTLCoder-DeepSeek-4bit shows a slight performance degradation compared to the original model, it is still superior to GPT-3.5 on the VerilogEval benchmark and comparable to it on RTLLM V1.1 with only 4GB size. Such RTLCoder-4bit can work on a simple laptop, allowing it to serve as a local assistant for engineers, addressing privacy concerns.
@@
As for the pre-trained model selection, we can see that different base model also has a significant impact on the performance of the fine-tuned model. On one hand, RTLCoder-DeepSeek slightly outperforms RTLCoder-Mistral in accuracy on most benchmarks. This trend is consistent with the base model’s relative accuracy (i.e., DeepSeek outperforms Mistral in most benchmarks). On the other hand, the inference speed of RTLCoder-Mistral is considerably faster than RTLCoder-DeepSeek, largely because of the Grouped Query Attention and Rolling Buffer KV Cache techniques used in Mistral.
IV-E Experiment Results in Detail
To further examine the performance in detail, for both benchmarks [11, 12], we report RTLCoder’s performance on each individual design case in both syntax and functionality correctness.
We list the test results of RTLCoder-Mistral and available baseline models on the RTLLM V1.1 benchmark for each design task in Table  IV. Given 5 trials of generation, here we counted the number of passed cases in terms of Syn-VCS, Syn-DC, and Functionality. As introduced, for both syntax and functionality, we count one success if any of the 5 trials pass the test. Generally, Syn-VCS is easier to pass than Syn-DC.
We further inspect the wrong answers in Table  IV. We observed that the overall code structures of wrong answers from GPT-3.5, GPT-4, and RTLCoder-Mistral exhibit no obvious mistakes, despite the functionality incorrectness. In comparison, the code generated by other open-source baselines occasionally contains obviously redundant content or deviates considerably from the given description. In terms of syntax, we observed that both GPT and RTLCoder-Mistral frequently assign 0 directly to two-dimensional arrays, resulting in syntax errors. Regarding functionality, we noticed that for more complex combinational logic circuits such as multi_pipe_4bit and multi_pipe_8bit, and sequential logic circuits like pulse_detect and FSM, some of the logical behaviors described in the instructions are not adequately captured by all LLM solutions, leading to functional errors.
The RTLCoder-Mistral’s results on VerilogEval Benchmark are reported in Figure 16. Each cell in the image represents one design case, with color indicating the number of successful ones among all 20 trails. There are 8 columns in each image. The location of cell $(i,j)$ represents the $((j-1)\times 8+i)^{\text{th}}$ design case in the provided description file. So we used white cells to fill the cells in the last row ($18^{\text{th}}$ row for EvalMachine and $20^{\text{th}}$ row in the EvalHuman) that do not correspond to a design task.
Table V: Ablation study of different decoding methods in RTLLM V1.1 Benchmark [11]. The result of the sampling decoding method is adopted and reported in the Table III.
Model Sampling decoding Beam search decoding
[used in experiment] [for ablation study]
Syn-VCS Syn-DC Func Syn-VCS Syn-DC Func
Thakur et al. [13] 86.2 86.2 24.1 69.0 51.7 17.2
StarCoder[22] 93.1 82.8 27.6 58.6 58.6 17.2
RTLCoder-Mistral-4bit 86.2 75.9 41.4 75.9 65.5 31.0
RTLCoder-Mistral 96.6 79.3 48.3 75.9 72.4 37.9
During the process of generating text sequences, the model continuously repeats the behavior of predicting the next token. For all models in our experiment, we adopt the sampling method, which randomly selects the next token from the vocabulary dictionary based on the probability distribution. Here we further add an ablation study based on the beam search method. A beam of the top “beam size” sub-sequences with the highest generation probabilities is maintained and updated during the generation process. We conduct experiments using beam search method with a beam size 5 on RTLLM V1.1 for RTLCoder-Mistral and open source baselines. The results are shown in Table  V. The accuracies of all methods drop after adopting beam search. RTLCoder-Mistral is still superior to all the open-source baselines with beam search.
Verilog Python Cpp Sh
Mistral-7B-v0.1 4.49 25.2 30.1 9.07
DeepSeek-Coder-6.7b-v1 30.2 66.8 63.9 36.3
RTL-Coder Mistral 36.7 25.3 21.4 1.99
RTL-Coder DeepSeek 41.6 66.7 63. 0 32.6
Table VI: The pass@1 results of the trained RTLCoder and two base models on different programming tasks
We also investigate the performance of the RTLCoder fine-tuned on the Verilog task compared to its pre-train models on other code-generation tasks. Table VI shows pass@1 results on Veilog tasks from VerilogEval-Human [12], python tasks from HumanEval+ [34], Cpp and Sh programming tasks from  [35]. RTLCoder-DeepSeek shows a significant improvement over its base model in the Verilog task and performs similarly to the base model in the other three programming tasks. RTLCoder-Mistral, compared to its base model, also shows better performance in the Verilog task but experiences more performance degradation in the Cpp and Sh benchmarks. This observation indicates that different pre-train models when fine-tuned on the same specific task, exhibit varying degrees of forgetting when applied to other benchmarks. Overall, such a degradation in other programming tasks does not affect our target application (i.e., circuit RTL generation).
V Limitation and Future Work
The results of GPT4 on VerilogEval-Human and RTLLM 1.1 are very impressive in Table III which indicates the significant superiority of GPT4 to all the other reported models. In our opinion, GPT-4, the most powerful large-scale model by OpenAI to date, outperforms existing open-source models for mainly three main reasons: 1) The size of the pre-training dataset used by GPT-4 is orders of magnitude larger than our adopted pre-trained models. The generation of a huge dataset that enables high-performance LLM consumes a significant amount of manpower and resources. 2) The GPT-4 model is much larger in scale than any existing open-source model. The scaling law  [36] suggests that as the model parameters increase, the model’s generalization and learning capabilities also improve. 3) The alignment technique such as RLHF [37] employed by GPT-4 can enhance the quality of the model’s outputs, reducing incorrect answers, which is also one of the motivations for us proposing the scoring-based training approach in our paper.
Given limited resources, we believe there are several key strategies for the open-source model to outperform GPT4 in specific tasks.
  • The diversity and coverage of the dataset can be further improved. Due to limited manual efforts during dataset generation, the training set primarily relies on guiding GPT to create instructions. To further enhance the coverage of the dataset for Verilog design problems, manual inspection and creation for high-quality data sample generation can be conducted.
  • The checking of functional correctness of the Verilog code should be further explored if possible, using either manpower or a reliable automated verification process. Although our fine-tuned models have a significant performance improvement and even outperform GPT-3.5, reliable filtering based on functionality checking can certainly further boost the model’s generation ability.
  • The automated functionality checking for hardware is challenging. A possible solution is to apply LLMs such as GPT to also help generate verification assertions based on instructions. Then the code combined with assertions can be verified by verification platforms such as Cadence Jasper. However, the correctness of assertions, since they are also automatically generated by LLMs, is not guaranteed either. This may result in wrongly filtering out many correct samples.
  • The scoring-based training scheme can be improved by applying automated functionality checking. This technique can hopefully enhance the model’s instruction alignment capability.
  • Leveraging RAG (Retrieval-Augmented Generation) technique [38] combined with high-quality datasets may improve the robustness of the model’s outputs.
VI Conclusion

### v5

@@ III New Training Scheme Incorporating Code Quality Feedback
The sequence generation is autoregressive, which means the model always predicts the next token based on its own generated previous ones rather than the reference tokens. Therefore, the traditional model tuning based on maximum likelihood estimation (MLE) would result in a phenomenon named exposure bias [23, 24] and the trained model would still generate many low-quality codes. To alleviate this phenomenon, we propose a new LLM training scheme that incorporates code quality scoring. It further improves the RTLCoder’s performance on the RTL generation task.
@@
For each instruction, we will now collect multiple additional code candidates generated by the initial pre-trained model. Then we pack these candidates and the original reference code $y_{i}$ together as $\textbf{y}_{i}=\left\{y_{i,k}\right\}$, $k=1,2,..,K$, where $K$ represents the number of generated code for one instruction ${x}_{i}$. Next, all these candidates will be scored by the scoring mechanism $R(x_{i},y_{i,k})$ which could be a syntax checker or unit test for functionality check. We will then obtain a set of score $\textbf{z}_{i}=\left\{z_{i,k}\right\}$, $k=1,2,..,K$, denoting the quality for the code sample $\{y_{i,k}\}$. In the training process, we make the model learn to assign relatively higher generation probabilities to answers with higher scores.
To further make this training scheme more memory efficient, we decompose the computation graph calculation and use the gradient accumulation-alike method to reduce the space complexity from $O(K)$ to $O(1)$.
@@ IV-A Evaluation Benchmark and Metric
To evaluate the performance of Verilog code generation, there are two representative benchmarks VerilogEval [12] and RTLLM [11]. The VerilogEval [12] benchmark consists of two parts, EvalMachine and EvalHuman, each including more than 100 RTL design tasks. We follow the original paper [12] and use the widely-adopted $pass@k$ metric. The RTLLM V1.1 [11] benchmark contains 29 RTL design tasks at a larger design scale. We use Synopsys VCS [25] to calculate the scores of the design syntax part and design functionality part separately. In both parts, following the original benchmark [11], each task is counted as success as long as any of 5 trials passes the test. This can be interpreted as pass@5 metric. In the generation process, we set ${top}_{p}=0.95$ and $temperature=\left\{0.2,0.5,0.8\right\}$. For all tested models (i.e., baselines, RTLCoder, and ablation studies), we evaluate all 3 temperaturetemperature conditions and report the best of each model.
@@
RTLCoder RTLCoder-Mistral 7B 62.5 72.2 76.6 36.7 45.5 49.2 96.6 48.3
@@
IV-B Model Training
To ensure a fair evaluation of our proposed RTLCoder, before training, we explicitly examined the similarity between samples in our proposed training dataset and those test cases in benchmarks [12, 11] using Rouge-L metric. Then we get rid of our training samples that are highly similar to test cases during the training process.
@@ IV-B Model Training
Based on our generated dataset with 27K instruction-code pairs, we choose the latest Mistral-7B-v0.1 [21] and DeepSeek-Coder-6.7b [26] as the basic pre-trained model for finetuning. In all experiments, we opted for the Adam optimizer with $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, and learning rate $\gamma$ = 1e-5, while abstaining from the use of weight decay. Concurrently, we established a context length of 2048 and a global batch size of 256. We trained the model on only 4 consumer-level RTX 4090 GPUs (24GB each), each of which could only afford $2\times 2048$ context length using DeepSpeed stage-2 [27].
@@
To implement our proposed training scheme, we first generated 3 code candidates for each instruction using a pre-trained model with Beam search method. Then we use Pyverilog [28] as the syntax checker to score the code candidates.
IV-C Experiment Results Overview
@@
Table II summarizes the comparison of all relevant RTL generation solutions, including commercial models GPT3.5/GPT4, models customized for Verilog generation [12, 13] [18], software code generators [19, 20, 21], and ablation studies of RTLCoder.
@@
In the VerilogEval benchmark [12], for both EvalMachine and EvalHuman categories, RTLCoder-DeepSeek scores 61.2 and 41.6 respectively. It clearly outperforms GPT-3.5 and is only inferior to GPT-4 among all the models in EvalHuman. Specifically, in the EvalMachine part, RTLCoder-DeepSeek and RTLCoder-Mistral even outperform GPT4 by an absolute value of 1.2% and 2.5%. A similar trend can be observed in the RTLLM benchmark V1.1 [11]. RTLCoder is second only to GPT-4. In summary, RTLCoder outperforms GPT-3.5 and all non-commercial baseline models in most of the metrics.
@@
Furthermore, we validate the effectiveness of our proposed dataset and algorithm through an ablation study. The RTLCoder-Mistral-Direct and RTLCoder-DeepSeek-Direct are directly trained with the traditional MLE method. Using our training dataset, they can already significantly outperform the base model and even GPT-3.5 on part of these indexes. Then the RTLCoders trained with our proposed training scheme further outperform those using Direct training method on all benchmarks, indicating that our training method greatly further improves the model performance.
@@
V Conclusion